in 3 years,alex will be 3 times as old as his sisters precy.a year ago,alex was 7 times as old as precy.how old are they now?
Answer by AAfter Search(61)

Let the ages of alex and precy be 'a' years and 'p' years.
Given, a + 3 = 3(p + 3) => a = 3p + 6 ... (i)
a - 1 = 7(p - 1) => a = 7p - 6 ... (ii)
From (i) and (ii), we get
3p + 6 = 7p - 6
=> 12 = 4p
=> p = 3
=> a = 3 x 3 + 6 = 9 + 6 = 15
Hence, the present ages of alex and precy are 15 years and 3 years.